OpenStudy (akgorgeous):

How many solutions does the following equation have? |6x + 12| = −18

OpenStudy (peachpi):

can absolute value be negative?

OpenStudy (alivejeremy):

No solutions.

OpenStudy (nevillesnetwork):

6x - 12 = 18 6x = 18 + 12 6x = 30 (divide both sides by 6 to get x) 6x/6 = 30/6 x = 5

OpenStudy (alivejeremy):

Your answer: No solutions.

OpenStudy (nevillesnetwork):

One solution.

OpenStudy (nevillesnetwork):

Wait wrong problem

OpenStudy (alivejeremy):

No.

OpenStudy (nevillesnetwork):

Let's solve your equation step-by-step. |6x+12|=−18 Solve Absolute Value. |6x+12|=−18 No solutions. (Absolute value cannot be less than 0.)

OpenStudy (alivejeremy):

bruh...........